OpenStudy (anonymous):

$100,000 for 20 years compounded at 4 percent annually results in a rate per period of:

OpenStudy (mathstudent55):

4%

OpenStudy (anonymous):

how did you get this answer

OpenStudy (mathstudent55):

It seems to me that the period is 1 year and the compounding is yearly too, so it's 4% per period.

OpenStudy (mathstudent55):

Maybe this means something else, and I misunderstood. The future value, F, of a present value, P, invested at i% for n compounding periods is: F = P(1 + i)^n In this case, F = 20,000(1 + 0.04)^20 F = 20,000(1.04)^20 F = 20,000(2.19112) F = 43,822.40 After having $20,000 at 4% per year compounded yearly, you end up with 43,822.40. That means you gained $43,822.40 - $20,000 = $23,822.40 in interest in 20 years. This means an average of $23,822.40/20 = $1191.12 per year. $1191.12 turns out to be 5.96% of $20,000. So it's an effective 5.95% simple interest rate per year.